OpenStudy (anonymous):

Aly is looking for a new car. She has test-driven two cars but can only purchase one. The probability that she will purchase car A is 0.47 and the probability that she will purchase car B is 0.35. What is the probability that she will not purchase either car A or car B? 0.12 0.41 0.82 0.18

OpenStudy (amistre64):

what is A + B ?

OpenStudy (amistre64):

if we assume that she will only purchase one car, then A+B + (other) = 1

OpenStudy (amistre64):

we know A and we know B ...

OpenStudy (anonymous):

0.47 + 0.3 = 0.82

OpenStudy (amistre64):

then: .82 +.12 ---- 1.00 so id say the probability of doing notA or notB is 1 - (A+B) = .12